# SpeciaLex: A Benchmark for In-Context Specialized Lexicon Learning

### v1

@@ Abstract
Specialized lexicons are collections of words with associated constraints such as special definitions, specific roles, and intended target audiences. These constraints are necessary for content generation and documentation tasks (e.g., writing technical manuals or children’s reading materials), where the goal is to reduce the ambiguity of text content and increase its overall readability for a specific group of audience. Understanding how large language models can capture these constraints can help researchers build better, more impactful tools for wider use beyond the NLP community. Towards this end, we introduce SpeciaLex, a benchmark for evaluating a language model’s ability to follow specialized lexicon-based constraints across $18$ diverse subtasks with $1,785$ test instances covering core tasks of Checking, Identification, Rewriting, and Open Generation. We present an empirical evaluation of $15$ open and closed-source LLMs and discuss insights on how factors such as model scale, openness, setup, and recency affect performance upon evaluating with the benchmark. [^1] We will release SpeciaLex including all research artifacts associated with its development (code, data, tasks) upon publication of this paper.
@@ 2 Related Work
Augmenting Lexicons and Dictionaries to LLMs. The use of lexicons and dictionaries has served as an additional knowledge base for LLMs across a number of tasks. He and Yiu (2022) used the Oxford dictionary to finetune BART models to generate appropriate sentence examples based on words. Yu et al. (2022) used dictionary definitions of rare words to improve the pre-training of LLMs. Similarly, Wu et al. (2022) also used specialized lexicons to improve the contrastive learning objective of pertaining BERT and RoBERTa models for tasks such as abusive language detection and sentiment analysis. Our use of lexicons for SpeciaLex serves as a reference of constraint for LLMs for content generation tasks.
@@ 3.2 Task Typology
T1 - Checking involves validation of a given input text whether to conforms to a specified constraint. As a validation task, the constraint can only be one of the three recognized SpeciaLex constraints. The outputs for Checking tasks are binary YES or NO.
T2 - Identification is another validation-type task that involves listing (non)conformity of an input text from a given task and lexicon-based constraint. The variation of Identification spans recognizing what word or set of words violate specific roles, special definitions, or target audience assigned by recognized constraints as well as identifying the most appropriate correct target audience.
T3 - Rewriting involves reconstructing an input text that violates a given lexicon-based constraint into a correct version which will be evaluated accordingly. We consider Rewriting as a semi-open generation task since the output is no longer structured like Checking or Identification, but the LLM still has a reference to the incorrect version and in-context demonstrations as guidance.
T4 - Open Generation is a full open-ended generative task that requires the LLM to generate a constraint-compliant output on-the-fly from the input text and task-specific demonstrations. Moreover, unlike Rewriting, each Open Generation task instance has no reference to an incorrect version and only the word and its associated constraint it needs to generate with, which makes this task more challenging.
@@ 4.2 Specialized Lexicon Data
For constructing the test cases in SpeciaLex, we use globally-recognized specialized lexicons in English, both used in technical writing and language assessment, to capture the three core constraints described in Section 3. Additional information can be found in Appendix C.
Simple Technical English Lexicon (STE) is an international industry-standard specification of controlled language used for simpler and clearer English technical documentation developed by the European Association of Aerospace Industries (AECMA). Previously exclusively used within aerospace engineering, STE has been adopted in many fields, including education, defense, and maintenance, and used across tasks such as machine translation and simplification Kuhn (2014). STE has a lexicon component that contains $1,259$ words with associated alternative words and part-of-speech information and $939$ with special definitions. These constraints aim to reduce ambiguity and ensure that the text can be easily understood by non-native English speakers. We use the lexicon of STE Issue 7 (released 2017) to manually construct test instances for the tasks classified evaluating Specific Roles and Special Definition constraints for SpeciaLex.
@@ 5.4 Error Analysis on Low-Performance Tasks
We take a closer look at the tasks with generally poor performances from models. This is particularly evident for tasks in Table 3 specifically on both Idenfitication subtasks requiring listing words from a text that are not recognized within the target audience level (ID13) and identifying the correct level (ID14). For the former, upon manual error analysis of model outputs, LLMs evaluated for the subtask often provide an insufficient number of required words (e.g., only giving $1-3$ words while the required is $5-6$), which includes words that are already within the recognized target audience level. For the latter, we see a trend where LLMs tend to oversimplify their estimations to lower levels (e.g., the correct level is B2, but models will give A2 or A1). We find similar insights from previous works on instruction-tuned LLMs oversimplifying level estimations for in-context learning tasks Imperial and Tayyar Madabushi (2023). We reserve the improvement of LLM performance for these specific subtasks for future work.
@@ 7 Conclusion
In this work, we introduced SpeciaLex, a benchmark for evaluating state-of-the-art LLMs in capturing specialized lexicon-based constraints for content generation tasks commonly prevalent across interdisciplinary areas such as education, technical writing, and engineering. We provided an in-depth and empirical exploration of model performance, including looking at the effects of model scale, openness, few-shot setup, and recency. Our findings support the use of open models such as Llama8-3B as good, competitive starting resources for the benchmark, which also serves as a step forward for accessible community adoption and springboarding to various domains.

### v2

@@ Abstract
Specialized lexicons are collections of words with associated constraints such as special definitions, specific roles, and intended target audiences. These constraints are necessary for content generation and documentation tasks (e.g., writing technical manuals or children’s reading materials), where the goal is to reduce the ambiguity of text content and increase its overall readability for a specific group of audience. Understanding how large language models can capture these constraints can help researchers build better, more impactful tools for wider use beyond the NLP community. Towards this end, we introduce SpeciaLex, a benchmark for evaluating a language model’s ability to follow specialized lexicon-based constraints across $18$ diverse subtasks with $1,785$ test instances covering core tasks of Checking, Identification, Rewriting, and Open Generation. We present an empirical evaluation of $15$ open and closed-source LLMs and discuss insights on how factors such as model scale, openness, setup, and recency affect performance upon evaluating with the benchmark. [^1] The task datasets and evaluation code can be found at: https://github.com/imperialite/specialex/.
@@ 2 Related Work
Augmenting Lexicons and Dictionaries to LLMs. The use of lexicons and dictionaries has served as an additional knowledge base for LLMs across a number of tasks. He and Yiu (2022) used the Oxford dictionary to finetune BART models to generate appropriate sentence examples based on words. Yu et al. (2022) used dictionary definitions of rare words to improve the pre-training of LLMs. Similarly, Wu et al. (2022) also used specialized lexicons to improve the contrastive learning objective of pertaining BERT and RoBERTa models for tasks such as abusive language detection and sentiment analysis. Our use of lexicons for SpeciaLex serves as a reference of constraint for LLMs for content generation tasks. Moreover, while all the previous works cited make use of extra training via finetuning to make their models task-specific, SpeciaLex focuses on capturing constraints purely by in-context learning while preserving the evaluated models’ ability to perform across general tasks.
@@ 3.2 Task Typology
T1 - Checking involves validation of a given input text whether to conforms to a specified constraint. As a validation task, the constraint can only be one of the three recognized SpeciaLex constraints. The outputs for Checking tasks are binary YES or NO. There are a total of $4$ Checking tasks in SpeciaLex.
T2 - Identification is another validation-type task that involves listing (non)conformity of an input text from a given task and lexicon-based constraint. The variation of Identification spans recognizing what word or set of words violate specific roles, special definitions, or target audience assigned by recognized constraints as well as identifying the most appropriate correct target audience. There are a total of $4$ Identification tasks in SpeciaLex.
T3 - Rewriting involves reconstructing an input text that violates a given lexicon-based constraint into a correct version which will be evaluated accordingly. We consider Rewriting as a semi-open generation task since the output is no longer structured like Checking or Identification, but the LLM still has a reference to the incorrect version and in-context demonstrations as guidance. There are a total of $5$ Rewriting tasks in SpeciaLex.
T4 - Open Generation is a full open-ended generative task that requires the LLM to generate a constraint-compliant output on-the-fly from the input text and task-specific demonstrations. Moreover, unlike Rewriting, each Open Generation task instance has no reference to an incorrect version and only the word and its associated constraint it needs to generate with, which makes this task more challenging. There are a total of $5$ Open Generation tasks in SpeciaLex.
@@ 4.2 Specialized Lexicon Data
For constructing the test cases in SpeciaLex, we use globally recognized specialized lexicons in English, both used in technical writing and language assessment described below, to capture the three core constraints described in Section 3. Note that these lexicons do not require any additional expert annotations as they are already off-the-shelf resources packaged as expert-developed datasets. Additional information can be found in Appendix C.
Simple Technical English Lexicon (STE) is an international industry-standard specification of controlled language used for simpler and clearer English technical documentation developed by the European Association of Aerospace Industries (AECMA). Previously exclusively used within aerospace engineering, STE has been adopted in many fields, including education, defense, and maintenance, and used across tasks such as machine translation and simplification Kuhn (2014); Zambrini and Chiarello (2023). STE has a lexicon component that contains $1,259$ words with associated alternative words and part-of-speech information and $939$ with special definitions. These constraints aim to reduce ambiguity and ensure that the text can be easily understood by non-native English speakers. We use the lexicon of STE Issue 7 (released 2017) to manually construct test instances for the tasks classified evaluating Specific Roles and Special Definition constraints for SpeciaLex.
@@ 5.4 Error Analysis on Low-Performance Tasks
We take a closer look at the tasks with generally poor performances from models. This is particularly evident for tasks in Table 3 specifically on both Idenfitication subtasks requiring listing words from a text that are not recognized within the target audience level (ID13) and identifying the correct target audience level (ID14). For the former, upon manual error analysis of model outputs, LLMs evaluated for the subtask often provide an insufficient number of required words (e.g., only giving $1-3$ words while the required is $5-6$), which includes words that are already within the recognized target audience level. For the latter, we see a trend where LLMs tend to oversimplify their estimations to lower levels (e.g., the correct level is B2, but models will give A2 or A1). We find similar insights from previous works on instruction-tuned LLMs oversimplifying level estimations for in-context learning tasks Imperial and Tayyar Madabushi (2023).
In hindsight, knowing that commercial and open LLMs may underperform for niche tasks that specific NLP tools or models can easily solve is something that domain-specific users should know when using these LLMs. Thus, we see this as an advantage that our benchmark exposes certain limitations that can inform the NLP community to build upon this work. We reserve the improvement of LLM performance for these specific subtasks for future work.
@@ 6 A SpeciaLex Guide
Is in-context learning better than domain-specific finetuning? ICL allows flexibility and preserves general model performance. The benefit of SpeciaLex by benchmarking via in-context learning is that it avoids re-training LLMs to one specific task only and preserves the user’s ability to use the LLM (or LLM interfaces such as ChatGPT) to perform other downstream tasks such question answering, summarizing content, and solving problems to name a few. Moreover, in case LLMs perform poorly on tasks via in-context learning, the results would serve as a helpful direction for domain users to then explore finetuning or other optimization methods. Nonetheless, this recommendation is not prescriptive if domain users’ priority is to develop a model that performs well in capturing constraints from only one source of reference via the specialized lexicon.
@@ 7 Conclusion
In this work, we introduced SpeciaLex, a benchmark for evaluating state-of-the-art LLMs in capturing specialized lexicon-based constraints for content generation tasks commonly prevalent across interdisciplinary areas such as education, technical writing, and engineering. We provided an in-depth and empirical exploration of model performance, including looking at the effects of model scale, openness, few-shot setup, and recency. Our findings support the use of open models such as Llama8-3B as good, competitive starting resources for the benchmark. In hindsight, SpeciaLex will serve as a reference guide for researchers within and outside of the NLP community, where they can check what specific models are good for certain task types (checking, identification, rewriting, generation) or handle specific constraints (specific roles, special definitions, target audience) and springboard further research based on their own domain specifications.
@@
Evaluation via In-Context Learning. In this work, we used prompting through in-context learning as one of the easiest ways users of various domain areas use an LLM (or LLM interfaces such as ChatGPT) with minimal effort. Moreover, the benefit of benchmarking via in-context learning is that it avoids fine-tuning or re-training the LLM to perform one specific task only and still preserves the LLM’s ability to perform language tasks such as summarizing, chatting, and answering questions. Lastly, evaluating through in-context learning can be considered the first step towards exposing the limitations of LLMs (as done by previous benchmarks) which can serve as a springboard for further domain-specific training or finetuning in future works.